A particle executes simple harmonic motion with an amplitude of 10 cm. At what distance from the mean position are the kinetic and potential energies equal?

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. . r = 10cm Because, K.E. = P.E. So (1/2) m ω2 (r2– y2) = (1/2) m ω2y2 r2 – y2 = y2 ⇒ 2y2 = r2 ⇒ y = r/√2 = 10/√2 = 5√2 cm form the mean position